# Understanding Parents’ Perceptions and Practices Toward Children’s Security and Privacy in Virtual Reality

### v1

@@ 4.3 Responsible Stakeholders and Desired S&P-Enhancing Features (RQ3)
We asked participants to describe the responsibilities of different stakeholders and features that they desired to enhance children’s S&P in VR. Our findings suggest: (1) Participants generally agreed a multi-stakeholder ecosystem must be established to make concerted efforts towards more S&P support for children using VR, with parents being the most accountable stakeholder. (2) While participants desired several features including real-time feedback and alerts with regard to children’s S&P in VR, they also had reservations about VR companies’ data surveillance.
A multisided VR ecosystem towards more S&P support for children. When naming perceived stakeholders and their responsibilities in mitigating S&P risks for children’s VR usage, all participants suggested more than one stakeholder. Taking all participants’ opinions together, we summarized four main stakeholders – parents, VR companies (including headsets and apps), governments, and schools/educators.
@@
Besides better knowing VR technology, another gap exists in parents’ understanding of the S&P practices across different VR products. We asked our participants if they had ever read any privacy policies in VR platforms, and almost all of them answered no, including those with more relevant expertise. This was in part due to the low readability of the policies, but also because of the inaccessibility of reading them through the headsets.
VR headset manufacturers and app companies (we combine them as “VR companies” in this paper) were also considered largely responsible for mitigating the S&P risks. In our interviews, a few participants expressed their concerns about the conflicts between users’ privacy and companies’ business models. In general, they believed that most of the existing S&P-enhancing efforts were primarily for promoting their public image. Correspondingly, prior work has also documented the trade-off that companies face between protecting children’s privacy and sustaining their business models [89, 90, 91]. To mitigate this concern, our participants suggested making S&P guidance part of a mandatory step before the use. As P18 suggested, “I’d prefer if it was in plain language with some short mandatory steps like watching a video… These steps should be very upfront.” Making privacy guidance shorter, clearer, mandatory, and upfront would be parents’ preferred practice, showing the company’s real effort in making the products more S&P-centered.
@@
Last, as a few participants pointed out, as VR becomes more and more prevalent, schools and educators also have the responsibility of teaching children about the nuances of data surveillance and self-disclosure in VR, compared to other technologies.
@@
As noted by participants, most of the existing moderation and defensive mechanisms in VR are reactive, e.g., blocking and reporting. However, going through the process of blocking and reporting was perceived as “worthless”, as it would not make up for the damage and would take too much time to receive feedback. As parents, our participants preferred to know immediately if harmful interactions were taking place. Participants who were concerned about bad influences also hoped to have red flags or alerts sent to them, e.g., an indicator of bad language. However, these ideas also raised some participants’ concerns, mainly for companies’ data surveillance. As P18 explained:
@@
According to participants, the same concern also applied to AI-based harm detection and secondary apps – all sounding helpful for parents to manage risks more effectively at the price of volunteering more sensitive data to companies.
A few participants proposed an alternative – sending parents reports with only basic but essential information. In doing so, participants were still able to get an insight into their children’s usage while not risking too much of their privacy. Since participants did not reach an agreement on how this “basic but essential” report mechanism should be designed in the interview, we will propose our recommendations in Section 5.
@@ 5 Discussion
In summary, our work has found an alarming status quo where parents of underage VR users generally demonstrate a lack of awareness regarding their children’s S&P in VR due to their conceptions of VR as still being “cutting-edge” and therefore “harmless” (RQ1). Despite parents’ lack of S&P awareness in VR, they still carry out a series of general rules toward at-home technology usage. Notably, when it comes to VR, parents primarily rely on active strategies as opposed to passive ones that pose significant technical constraints (RQ2). Finally, by prompting our participants in the interviews, they collectively demonstrate a need for a multisided VR S&P-enhancing ecosystem, with an emphasis on parents’ responsibilities and expectations toward a series of S&P-enhancing features (RQ3). In the remainder of this section, we first discuss how our work extends or contradicts prior work. With the takeaways, we then discuss our recommendations for the critical stakeholders, including parents, educators, VR companies, and governments.
@@
Prior work has pointed out that parents’ incomplete understanding of technology compared to children might impede parents’ ability to engage in children’s technology use [23, 24]. As a result, parents might underestimate the amount and type of technology use and online interaction by their children, hence putting them at risk [92, 69, 68]. Similarly, while bystanders in social VR have witnessed great risks for underage users [4], we noticed that parents in our study, especially those with fewer VR experiences, claimed that their perceived S&P risks would not apply to their own children at the current stage (Section 4.1.1). Therefore, it is worth suspecting whether the parents had underestimated children’s VR activity and S&P risks. Considering our focus on parents’ S&P implications in VR, we did not interview the children of our participants. This observation should be further studied by involving child participants in future work.
Parents’ overgeneralization of S&P rules in all digital devices defer addressing concerns to the future. An important finding of our work is that parents overgeneralize their S&P rules across diverse contexts, including sharing the same set of rules for all forms of digital communications (Section 4.2.1). Despite the fact that parents understand the potential of extensive biometric data being collected in VR compared to other technologies [37, 13], most of them still discard the differences in practice, especially those who seldom or never use internet-enabled VR applications. We believe this phenomenon may be attributed to parents’ lack of experience in using VR and their knowledge gap in the nuances between VR and other technologies. We will propose relevant recommendations for parents and educators later in this section.
Another reason for this type of overgeneralization is that parents prioritize children’s S&P in other everyday technologies such as social media (Section 4.1.1). For this reason, parents deferred addressing their S&P concerns in VR. Interestingly, in a study of children’s S&P online in general (e.g., social media) [16], parents still consider S&P risks as a future concern. Therefore, we suggest that in addition to enhancing parents’ understanding of VR technologies, there should be other measures to enhance S&P education itself more effectively, such as setting physical references as accounted below.
@@
Problematically, while most parents understand the differences between S&P risks in offline settings and VR, the proximity in social interaction between VR and face-to-face settings has been overlooked. Only P5 and P7 who have substantial VR experiences educate their children about S&P risks in VR by setting physical references or metaphors (Section 4.2.1). All the other parents in our study primarily rely on their rules for other technologies, e.g., not disclosing privacy when chatting with people online. However, text-based communication is drastically different from voice chat in VR [43], which resembles talking to friends in face-to-face scenarios. This shows a need to set physical references or metaphors when educating children about S&P in VR. As suggested by prior work [93, 94, 95, 96], physical metaphors are important in improving S&P awareness and protective strategies. For example, food nutrition labels serve as a physical metaphor for “privacy nutrition labels”, better informing users of S&P-related information [82, 97, 98].
Parents’ complex tradeoffs between passive and active strategies raised questions toward S&P controls in VR. Our work extends prior work that discussed children’s autonomy and S&P in their technology use [92, 22, 99, 16], suggesting that parents navigate complex tradeoffs between children’s autonomy and S&P in VR (Section 4.2.3). For example, even though parents are largely concerned about children’s self-disclosure in social interaction, they consider parental controls that restrict social communications to be less useful because such an “all-or-nothing” approach would not help children learn the correct tactics in practice (Section 4.2.2).
Notably, one difference between VR and other everyday technologies lies in the fact that VR has not yet “gotten to the point where having a phone and a tablet is necessary for you to be on par with your peers,” as highlighted by P1. Due to this perceived difference, some parents consider maintaining children’s S&P to be much more important than their autonomy in VR.
@@
For parents, in particular, it is recommended to experience VR themselves as much as possible. It would be even better if parents could maintain their knowledge level about VR close to or even beyond their children’s. Since VR is an immersive technology, the social experience and data collection might be completely different from what they know about smartphones, trying out VR themselves would be the best way for parents to understand the nuances. The significance of parents’ firsthand experience is also obvious in our findings, as participants with more VR experiences demonstrated better awareness and richer insights.
The same recommendation also applies to educators. As school education has already incorporated S&P risks in social media as part of the curriculum [23], it is also crucial to list VR S&P risks in the agenda, especially for those who utilize VR as an education technology [100, 101, 102, 12].
@@
To improve the usability of S&P controls in VR, two main problems must be addressed – (1) technical constraints such as connection issues and high latency, and (2) low engagement. Among a series of controls, parents find headset mirroring most effective (Section 4.2.2). However, due to technical issues, they also refrain from using this feature. In place of headset mirroring, parents hope to use secondary apps accompanied by the VR product (Section 4.3). The secondary app can not only just replace the functionality of headset mirroring, but also provide real-time alerts and other features to help parents gain more engagement while protecting their children in a more timely manner.
Considering parents’ S&P concerns about these desired features, we are proposing to also increase the granularity of S&P controls. Specifically, since parents all have very different considerations when facing the tradeoffs between children’s S&P from the company and strangers, S&P controls should enable parents to set what exactly information they allow companies to collect, analyze, and send real-time or periodic S&P alerts to parents.
In addition, almost all the parents in our study have never read the privacy policies of VR products, showing an urgent need to improve the S&P guidance for parents and children in VR. In short, S&P guidance in VR should be modality-adapted, mandatory, and engaging. Indeed, reading lengthy policies in VR can be difficult [103], and parents seldom search the S&P practices of VR products using mobile phones or PCs. Therefore, it is necessary to adapt the design of texts to best align with users’ reading preferences in VR, considering text lengths, sizes, color contrasts, etc.  [103].
Besides valuing users’ reading experiences in VR, other mechanisms include having an interactive onboarding tutorial to introduce potential S&P-risky scenarios and protective mechanisms. In particular, as prior work suggests [5], a “straightforward and well-explained guideline” in line with minors’ literacy levels is important.
@@ 6 Conclusion
This study presents an in-depth qualitative analysis of parents’ perceptions and practices toward children’s S&P in VR. Our results highlight that parents generally lack or discard S&P awareness in children’s use of VR. In mediating children’s use of VR, parents currently primarily rely on active strategies such as verbal education. Passive strategies, however, have been less effective due to technical constraints. Parents also agree that multiple stakeholders are responsible for mitigating S&P risks for children, with parents being the most accountable stakeholders. To conclude, we propose specific and actionable recommendations for the critical stakeholders.
@@ Acknowledgments
We thank Sparsha Perupalli and anonymous reviewers for their constructive feedback. We also thank Iffat Anjum and Nazanin Sabri for helping with our pilot interviews. We further thank our participants for their insightful input. This material is based upon work supported in parts by a Meta Research gift award. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Meta Inc.
Appendix A Interview Protocol
A.1 Introduction
Hi there! Thank you for participating in our study! My name is [NAME] and this is [NAME]. In this interview, I will be asking you about your usage of VR and your children’s usage of VR.
Before we get started, do you have any questions about the consent form that you signed? [If yes, answer the questions].
Before we begin the interview session, can you please show me your VR headset(s) to the camera? [If they decline, politely convey that we cannot continue the interview process]. As mentioned in the consent form, we would like to audio-record this interview for analysis purpose. We will keep our video on throughout the interview process; but it is completely up to you if you would like to have your video on or off. Could you please confirm that you allow us to audio-record this interview? [Start the audio recording].
We will use the term “children” in general, but if your response would be different for any specific child of yours, please feel free to let us know. We are very interested in understanding this. Thank you again. Let’s get started!
A.2 VR Usage
Parent’s Use of VR
  • Do you use VR?
  • Could you please tell us the purposes for which you use VR?
  • Could you please name the VR apps that you use?
  • How long have you been using VR and how often do you use VR?
Children’s Use of VR
  • Does any of your children under the age of 18 use VR?
  • Could you please tell us the purposes for which your children use VR?
  • Could you please name the VR apps that your children use?
  • Since what age have your children started using VR, and how often do they use VR?
A.3 Parents’ Perceived Risks of Children’s Use of VR in General
  • What group or demographic do your children usually interact with on VR? In terms of their age, gender, race, education, or any other demographic information that you have.
  • From your perspective, what are the potential risks to (your) children because of using VR?
  • What has made you perceive these potential risks to (your) children’s use of VR? How / in what way?
  • Who do you think is responsible for mitigating the potential risks? Please name any individuals, groups, or institutions that you perceive to be responsible for mitigating the risks for children.
  • What are these stakeholders’ responsibilities in mitigating the potential risks?
A.4 Parents’ Perceived Concerns Toward Their Children’s S&P in VR
  • On a scale of 1 to 5, one being not at all concerned and five being very concerned, how concerned about security and privacy regarding your children’s use of VR?
    • If “Not at all concerned”:
      What makes you not at all concerned about your children’s use of VR?
    • If concerned:
      What makes you concerned about your children’s use of VR?
  • Have you ever had any privacy or security concerns about (your) children’s use of VR?
    • If yes:
      What privacy or security concerns do you have about (your) children’s use of VR?
    • If no:
      What privacy or security risks do you imagine there would be?
A.5 Parents’ S&P Protective Practices Toward Their Children’s Use of VR
  • Have you ever taken any steps to address your concerns about your children’s use of VR? If yes, please tell us what actions you have taken and why? (ex: do you monitor them)
    • How do your children react / respond to these steps?
  • Have you ever read about the privacy practices of the VR platform that (your) children interact with? The platform includes VR apps as well as a VR headset. Where did you get this information?
    • Have you been satisfied with the information you found? What additional information about the privacy and security practices of the VR platform would you like to know about?
    • How would you like this information to be presented to you?
  • Have your children ever talked to you about their experience with using VR? What was the conversation about?
  • Have you ever discussed the privacy risks of VR with your children?
    If yes:
    • Could you please tell us about such conversations? For example, what topics did you talk about, and how did your children perceive the conversation?
    • Did you observe any changes in your children’s interaction with VR after such conversations, and if yes, in what way?
    If yes/no:
    • What is your perspective on the impact of having such a conversation with your children?
That brings us to the end of our session today! To record your completion of the interview, can you please fill out the compensation survey? (share link)
A.6 Live survey about parental control features
Parental Control Features
  • Headset mirroring
  • Set passwords
  • Block content
  • Block apps
  • Restrict playtime
  • Restrict social communication
For each feature, we asked parents two questions as follows:
  • Have you heard of this feature?
    • Yes
    • No
  • How effective do you think this feature is?
    • Not effective at all
    • Slightly effective
    • Moderately effective
    • Very effective
    • Extremely effective
Appendix B Parent Participants’ and Their Child(ren)’s VR Usage
TABLE II: Parents’ VR usage.
ID VR Device Ownership Parent Purposes of Usage Others
P1 Meta Quest Pro Creativity, Education, Gaming, Health, Productivity, Socialization, Streaming
P2 Meta Gaming, Other (please specify) Work
P3 Meta Quest 2 Gaming, Other (please specify) Exploration of VR
P4 Oculus Creativity, Education, Gaming, Productivity, Socialization
P5 Meta Quest 2 Gaming, Health, Socialization, Streaming
P6 Meta Quest 2 Creativity, Gaming, Socialization
P7 Meta Quest Gaming
P8 Meta Oculus Quest 2 Creativity, Education, Gaming, Health, Streaming, Other (please specify)
P9 DEVASCO & Bnext Creativity, Education, Health, Other (please specify) Virtual tours
P10 Oculus, Google Cardboard Education
P11 Oculus, Microsoft Hololens Creativity, Education, Gaming, Productivity, Other (please specify) Product training
P12 Oculus Creativity, Gaming, Socialization, Streaming
P13 Oculus Gaming, Health
P14 Oculus Meta Quest Gaming
P15 Oculus Gaming, Socialization
P16 Oculus Quest 2 Education, Streaming
P17 Meta Oculus Education, Streaming
P18 Oculus Gaming
P19 Oculus Quest 2 Education, Gaming, Health, Socialization, Streaming
P20 Oculus Creativity, Gaming, Socialization, Streaming
TABLE III: Children’s VR usage.
ID Child Purposes of Usage Others
P1 Unspecified
P2 Other (please specify) No longer interested
P3 Gaming
P4 Creativity, Education, Gaming
P5 Gaming
P6 Other (please specify) No longer interested
P7 Gaming
P8 Creativity, Education, Gaming, Socialization
P9 Creativity, Education, Health, Other (please specify) Virtual field trips
P10 Education, Gaming
P11 Gaming
P12 Unspecified
P13 Gaming, Health
P14 Gaming
P15 Gaming, Socialization
P16 Gaming
P17 Gaming, Streaming
P18 Gaming
P19 Gaming
P20 Gaming
Appendix C Interview Participants’ Full Demographics
TABLE IV: Parents’ age, gender, race, and education background distributions.
Age Gender Race Education Background
25-34 10% M 70% White 55% Doctorate degree 5%
35-44 15% F 30% Black 20% Master’s degree 45%
45-54 65% Asian 20% Professional degree 5%
55-64 10% Prefer not to say 5% Bachelor’s degree 30%
Associate’s degree 5%
1 or more years of college credit, no degree 5%

### v3

@@ 4.3 Responsible Stakeholders and Desired S&P-Enhancing Features (RQ3)
We asked participants to describe the responsibilities of different stakeholders and features that they desired to enhance children’s S&P in VR. Our findings suggest: (1) Participants generally agree a multi-stakeholder ecosystem must be established to make concerted efforts towards more S&P support for children using VR, with parents being the most accountable stakeholder. (2) While participants desire several features including real-time feedback and alerts with regard to children’s S&P in VR, they also have reservations about VR companies’ data surveillance.
A multisided VR ecosystem towards more S&P support for children. When naming perceived stakeholders and their responsibilities in mitigating S&P risks for children’s VR usage, all participants suggested more than one stakeholder. Taking all participants’ opinions together, we summarize four main stakeholders – parents, VR companies (including headsets and apps), governments, and schools/educators.
@@
Another gap exists in parents’ understanding of the S&P practices and policies across different VR products. We asked our participants if they had ever read any privacy policies on VR platforms, and almost all of them answered no, including those with more relevant expertise. This was in part due to the low readability of the policies, but also because of the inaccessibility of reading them through the headsets.
VR headset manufacturers and app companies (we combine them into “VR companies” in this paper) were also considered largely responsible for mitigating the S&P risks. In our interviews, a few participants expressed their concerns about the conflicts between users’ privacy and companies’ business models. In general, they believed that most of the existing S&P-enhancing efforts were primarily for promoting their public image. Correspondingly, prior work has also documented the trade-off that companies face between protecting children’s privacy and sustaining their business models [96, 97, 98]. To mitigate this concern, our participants suggested making S&P guidance part of a mandatory step before the use. As P18 suggested, “I’d prefer if it was in plain language with some short mandatory steps like watching a video… These steps should be very upfront.” Making privacy guidance shorter, clearer, mandatory, and upfront would be parents’ preferred practice, showing the company’s real effort in making the products more S&P-centered.
@@
Last, as a few participants pointed out, as VR becomes more and more prevalent, schools and educators also have the responsibility in teaching children about the nuances of data surveillance and self-disclosure in VR, compared to other technologies.
@@
As noted by participants, most of the existing moderation and defensive mechanisms in VR are reactive, e.g., blocking and reporting. However, going through the process of blocking and reporting was perceived as “worthless,” as it would not make up for the damage and would take too much time to receive feedback. As parents, our participants preferred to know immediately if harmful interactions were taking place. Participants who were concerned about bad influences also hoped to have red flags or alerts sent to them, e.g., an indicator of bad language. However, these ideas also raised some other participants’ concerns, mainly for companies’ data surveillance. As P18 explained:
@@
According to participants, the same concern also applied to AI-based harm detection and secondary apps, all sounding helpful for parents in managing risks more effectively at the price of volunteering more sensitive data to companies.
A few participants proposed an alternative – sending parents reports with only basic but essential information. In doing so, participants were still able to get an insight into their children’s usage while not risking too much of their privacy. As participants did not reach an agreement on how this “basic but essential” report mechanism should be designed in the interview, we will further propose our recommendations in Section 5.
@@ 5 Discussion
In summary, our work has found an alarming status quo where parents of underage VR users generally demonstrate a lack of awareness regarding their children’s S&P in VR due to their conception of VR being “cutting-edge” and therefore “harmless” (RQ1). Despite parents’ lack of S&P awareness in VR, they carry out a series of general rules toward at-home technology use. Notably, when it comes to VR, parents primarily rely on active strategies as opposed to passive ones that pose significant technical constraints (RQ2). Finally, by prompting our participants in the interviews, they collectively demonstrate a need for a multisided VR S&P-enhancing ecosystem, with an emphasis on parents’ responsibilities and expectations toward a series of S&P-enhancing features (RQ3). In the remainder of this section, we first discuss how our work extends or contradicts prior work. With the takeaways, we then discuss our recommendations for the critical stakeholders, including parents, educators, VR companies, and governments.
@@
Prior work has pointed out that parents’ incomplete understanding of technology compared to children might impede parents’ ability to engage in children’s technology use [23, 24]. As a result, parents might underestimate the amount and type of technology use and online interactions that their children engage in, hence putting them at risk [99, 71, 70]. Similarly, while bystanders in social VR have witnessed great risks for underage users [4], we noticed that parents in our study, especially those with less VR experience, claimed that their perceived S&P risks would not apply to their own children at the current stage (Section 4.1.1). Therefore, it is worth suspecting whether the parents had underestimated children’s VR activity and S&P risks. Admittedly, parents’ perception of VR’s immaturity may be correct in some aspects, especially regarding the immature hardware adversaries [40]. Most parents without much S&P experience overlooked or disregarded server adversaries that inferred sensitive data [40]. In other words, while parents may be correct that immersive data collection had not yet been quite extensive, data inference might have been largely underestimated, according to the established threat model [40].
Parents’ overgeneralization of S&P rules in all digital devices defer addressing concerns to the future. An important finding of our work is that parents overgeneralize their S&P rules across diverse contexts, including sharing the same set of rules for all forms of digital communication (Section 4.2.1). Despite the fact that parents understand the potential of extensive biometric data being collected in VR compared to other technologies [37, 13], most of them still discard the differences in practice, especially those who seldom or never use internet-enabled VR applications. We believe this phenomenon may be attributed to parents’ lack of experience in using VR and their knowledge gap in the nuances between VR and other technologies. We will propose relevant recommendations for parents and educators later in this section.
Another reason for this type of overgeneralization is that parents prioritize children’s S&P in other technologies such as social media (Section 4.1.1). For this reason, parents deferred addressing their S&P concerns in VR. Interestingly, in a study of children’s S&P online in general (e.g., social media) [16], parents still consider S&P risks as a future concern. Therefore, we suggest that in addition to enhancing parents’ understanding of VR technologies, there should be other measures to enhance S&P education itself more effectively, such as setting physical references as accounted below.
@@
Problematically, while most parents understand the differences between S&P risks in offline settings and VR, the proximity in social interaction between VR and face-to-face settings has been overlooked. Only P5 and P7 who have abundant VR experience educate their children about S&P risks in VR by setting physical references or metaphors (Section 4.2.1). All the other parents in our study primarily rely on their general rules for technology use, e.g., not disclosing privacy when chatting with people online. However, text-based communication is drastically different from voice chat in VR [45], which resembles talking to friends in face-to-face scenarios. This shows a need to set physical references or metaphors when educating children about S&P in VR. As suggested by prior work [100, 101, 102, 103], physical metaphors are important in improving S&P awareness and protective strategies. For example, food nutrition labels serve as a physical metaphor for “privacy nutrition labels,” better informing users of S&P-related information [89, 104, 105].
Parents’ complex tradeoffs between passive and active strategies raised questions toward S&P controls in VR. Our work extends prior work that has discussed children’s autonomy and S&P in their technology use [99, 22, 106, 16], suggesting that parents navigate complex tradeoffs between children’s autonomy and S&P in VR (Section 4.2.3). For example, even though parents are largely concerned about children’s self-disclosure in social interaction, they consider parental controls that restrict social communication less useful because such an “all-or-nothing” approach would not help children learn the correct tactics in practice (Section 4.2.2).
Notably, one difference between VR and other consumer-facing technologies lies in the fact that VR has not yet “gotten to the point where having a phone and a tablet is necessary for you to be on par with your peers,” as highlighted by P1. Due to this perceived difference, some parents consider maintaining children’s S&P much more important than their autonomy in VR. For these parents, passive strategies such as using parental controls are expected to be helpful.
@@
For parents, in particular, it is recommended to experience VR themselves as much as possible. It would be even better if parents could maintain their knowledge level about VR on par with their children’s. Since VR is an immersive technology, the social experience and data surveillance might be completely different from what they know about smartphones, trying out VR themselves would be the best way for parents to understand the nuances. The significance of parents’ firsthand experience in VR is also obvious in our findings, as participants with more VR experience demonstrated better awareness and richer insights.
A specific example is that parents with S&P experience demonstrated greater concerns toward hardware and server adversaries [40]. While the lay parents’ perception of immature VR technologies may be correct in terms of data collection, their ignorance of the risks of data inference in VR is alarming. Hence, we also urge more privacy education on the different types of adversary types in VR, especially server adversaries.
The same recommendations also apply to educators. As school education has already incorporated S&P risks in social media as part of the curriculum [23], it is also crucial to list VR S&P risks in the agenda, especially for those who utilize VR as an education technology [107, 108, 109, 12].
@@
In improving the usability of S&P controls in VR, two main problems must be addressed – (1) technical constraints such as connection issues and high latency, and (2) low engagement. Among a series of controls, parents find headset mirroring most effective (Section 4.2.2). However, due to technical issues, they also refrain from using this feature. In place of headset mirroring, parents hope to use secondary apps accompanied by the VR product (Section 4.3). The secondary app can not only replace the functionality of headset mirroring but also provide real-time alerts and other features to help parents gain more engagement while protecting their children in a more timely manner.
Considering parents’ S&P concerns about these desired features, we suggest increasing the granularity of S&P controls. Specifically, since parents all have very different considerations when facing the tradeoffs between children’s S&P from the company and strangers, S&P controls should enable parents to set what exact information they allow companies to collect and analyze for generating real-time or periodic S&P alerts.
In addition, almost all the parents in our study have never read any privacy policies of VR products, showing an urgent need to improve the S&P guidance for parents and children in VR. In short, S&P guidance in VR should be modality-adapted, mandatory, and engaging. Indeed, reading lengthy policies in VR can be difficult [110], and parents seldom search the S&P practices of VR products using mobile phones or PCs. Therefore, it is necessary to adjust the design of texts to align with users’ reading preferences in VR, considering text lengths, sizes, color contrasts, etc.  [110].
Besides valuing users’ reading experiences in VR, other mechanisms include interactive onboarding tutorials to introduce S&P-risky scenarios and protective mechanisms. In particular, as prior work suggests [5], a “straightforward and well-explained guideline” in line with minors’ literacy levels is important.
@@ 6 Conclusion
This study presents an in-depth qualitative analysis of parents’ perceptions and practices toward children’s S&P in VR. Our results highlight that parents generally lack or discard S&P awareness in children’s VR usage. In mediating children’s VR usage, parents currently primarily rely on active strategies such as verbal education. Passive strategies, however, have been less effective due to technical constraints. Parents also agree that multiple stakeholders are responsible for mitigating S&P risks for children, with parents being the most accountable stakeholder. To conclude, we propose specific and actionable recommendations for the critical stakeholders.
@@ Acknowledgments
We thank Sparsha Perupalli, anonymous reviewers, and the shepherd for their constructive feedback. We also thank Iffat Anjum and Nazanin Sabri for helping with our pilot interviews. We further thank our participants for their insightful input. This material is based upon work supported in parts by a Meta Research gift award. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Meta Inc.
Appendix A Usage and Demographic Tables
TABLE II: Parents’ and Children’s VR usage.
ID VR Device Ownership Parent’s VR Usage Children’s VR usage.
P1 Meta Quest Pro Creativity, Education, Gaming, Health, Productivity, Socialization, Streaming Unspecified
P2 Meta Gaming, Work No longer interested
P3 Meta Quest 2 Gaming, Exploration of VR Gaming
P4 Oculus Creativity, Education, Gaming, Productivity, Socialization Creativity, Education, Gaming
P5 Meta Quest 2 Gaming, Health, Socialization, Streaming Gaming
P6 Meta Quest 2 Creativity, Gaming, Socialization No longer interested
P7 Meta Quest Gaming Gaming
P8 Meta Oculus Quest 2 Creativity, Education, Gaming, Health, Streaming Creativity, Education, Gaming, Socialization
P9 DEVASCO & Bnext Creativity, Education, Health, Virtual Tours Creativity, Education, Health, Virtual Tours
P10 Oculus, Google Cardboard Education Education, Gaming
P11 Oculus, Microsoft Hololens Creativity, Education, Gaming, Productivity, Work Gaming
P12 Oculus Creativity, Gaming, Socialization, Streaming Unspecified
P13 Oculus Gaming, Health Gaming, Health
P14 Oculus Meta Quest Gaming Gaming
P15 Oculus Gaming, Socialization Gaming, Socialization
P16 Oculus Quest 2 Education, Streaming Gaming
P17 Meta Oculus Education, Streaming Gaming, Streaming
P18 Oculus Gaming Gaming
P19 Oculus Quest 2 Education, Gaming, Health, Socialization, Streaming Gaming
P20 Oculus Creativity, Gaming, Socialization, Streaming Gaming
TABLE III: Parents’ age, gender, race, education background, and recruitment channel distributions.
Age Gender Race Education Background Recruitment Channel
25-34 10% M 70% White 55% Doctorate degree 5% Personal connections 5%
35-44 15% F 30% Black 20% Master’s degree 45% LinkedIn 10%
45-54 65% Asian 20% Professional degree 5% Facebook 10%
55-64 10% Prefer not to say 5% Bachelor’s degree 30% Discord 10%
Associate’s degree 5% Reddit 5%
1 or more years of college credit, no degree 5% School Advertisement 60%